A 3.20 L sample of neon at 4.57 atm is added to a 11.0 L cylinder that contains argon. If the pressure in the cylinder is 5.13 atm after the neon is added, what was the original pressure (in atm) of argon in the cylinder? I keep on gettting 5.29
Answer #1

By PU, = P2 W2 (3:20) (4:57) = P2 (110) P2 = 1.329 atm. PAR = 5.13- Pie = 5-13 - 1.329 Par = 3.80 atm